OpenStudy (anonymous):

3/5f+24=4-1/5f (fractions, not division) Please show work.

OpenStudy (anonymous):

Just to be clear: \[\frac{3}{5}f + 24 = 4 - \frac{1}{5}f\]What have you tried so far?

jhonyy9 (jhonyy9):

(3f +120)/5 = (20- f)/5 (3f+120-20+f)/5 = 0 (4f+100)/5 =0 4f +100 =0 4f= -100 f= -100/4 f= -25